OpenStudy (anonymous):

The center of a circle is at (4, -1) and it has a radius of 6. What is the equation of the circle?

OpenStudy (anonymous):

Answers: A.) (x-4)^2 + (y + 1)^2 = 6 B.) (x-4)^2 + (y + 1)^2 = 36 C.) (x + 4)^2 + (y-1)^2 = 6 D.) (x + 4)^2 + (y-1)^2 = 36 I just want to know which one it is please :)

OpenStudy (marissalovescats):

Equation of a circle: